Tetrahymena thermophila protozoa have a minimum doubling time of 6.5 hours when grown using bacteria as the limiting substrate. The yield of protozoal biomass is 0.33 g per g of bacteria and the substrate constant is 12 mg/ L. The protozoa are cultured at steady state in a chemostat using a feed stream containing 10 g 1 of nonviable bacteria. What is the maximum dilution rate for operation of the chemostat?
1
Expert's answer
2020-09-01T07:12:00-0400

Doubling time is given as ,

Td= 6.5 , We know that

Dilution rate= "\frac{F}{V}" , here F=Volumetric flow rate, V= Volume

and under case of limiting substrate

"\mu_{max}= \frac{0.693}{T_d}"

And in case of limiting substrate, "\mu_{max}=D" here, D= Dilution rate

so,

"D=\frac{0.693}{T_d}"


"D=\frac{0.693}{6.5}"


D= 0.1066 "hr^{-1}" , This is dilution rate